# Taking the Guesswork Out of Language Model Evaluations

### v1

@@ Abstract
Language model users often issue queries that lack specification, where the context under which a query was issued—such as the user’s identity, the query’s intent, and the criteria for a response to be useful—is not explicit. For instance, a good response to a subjective query like “What book should I read next?” would depend on the user’s preferences, and a good response to an open-ended query like “How do antibiotics work against bacteria?” would depend on the user’s expertise. This makes evaluation of responses to such queries an ill-posed task, as evaluators may make arbitrary judgments about the response quality. To remedy this, we present contextualized evaluations, a protocol that synthetically constructs context surrounding an underspecified query and provides it during evaluation. We find that the presence of context can 1) alter conclusions drawn from evaluation, even flipping win rates between model pairs, 2) nudge evaluators to make fewer judgments based on surface-level criteria, like style, and 3) provide new insights about model behavior across diverse contexts. Specifically, our procedure uncovers an implicit bias towards WEIRD contexts in models’ “default” responses and we find that models are not equally sensitive to following different contexts, even when they are provided in prompts.
@@ 1 Introduction
Users of language models often issue queries that are underspecified Spärck-Jones et al. (2007); Clarke et al. (2009); Ziegler et al. (2019); Keyvan and Huang (2022); Herlihy et al. (2024), but common evaluation practices for language models do not account for this. Consider an evaluator presented with a language model’s response to an underspecified query such as “Is coffee good for you?” (Figure 1). A language model might respond with an explanation about benefits like antioxidants and mental alertness, but this output would be unacceptable to users with certain health conditions. Can an evaluator make an informed judgment about response quality without further details about the intended user’s preferences, background or criteria for the response to be useful?
@@
Table 1: Queries from five LM benchmark datasets, categorized based on the amount of underspecification. Queries can present multiple types (e.g., “What is the best team in the league?” is both incomplete and subjective).
In this work, we consider the role of context in the evaluation of language model responses to underspecified queries. We propose contextualized evaluations, a protocol to synthetically generate and incorporate diverse contexts (represented as follow-up question-answer pairs) to underspecified queries. By applying this procedure to queries from widely-used language model benchmark datasets, we investigate three main research questions:
First, we investigate whether providing context to evaluators has a substantial effect on the conclusions drawn from evaluation. We sample responses from pairs of language models and collect pairwise preference judgments from both model-based as well as human evaluators, in context-agnostic (only the query and model outputs) and context-aware settings (additionally with follow-up questions and answers to clarify the query). Our experiments show not only that inclusion of context during evaluation can significantly improve agreement between evaluators (3-10% absolute), but also that context-aware evaluation can even flip win rates between model pairs. This raises concern for the reliability of findings produced from context-agnostic evaluations using today’s language model benchmarks, which we find to be full of underspecified queries (Table 1).
Finally, we investigate whether context helps us learn more about the ability of models to adapt to different user contexts. In our work, we use contexts to study biases exhibited by “default” model responses, i.e., those generated without context (§5), to underspecified queries. We find strong evidence of a bias where default model responses are better aligned with WEIRD (Western, Educated, Industrialized, Rich and Democratic) contexts Henrich et al. (2010). Context can also allow us to directly evaluate the instruction-following and personalization abilities of models. We show this by evaluating sensitivity of models to different contextual attributes, where we find considerable disparity in the ability of models to adapt to different contexts, even when they’re provided in prompts6).
Figure 2: Types of common contextual attributes which can be lacking in underspecified queries, along with examples of follow-up questions for each attribute.
In summary, our findings suggest that underspecification can have a significant impact on the conclusions and insights drawn from evaluation. To address this, we propose contextualized evaluations, a simple and broadly applicable solution that involves synthesizing relevant context and injecting it into existing evaluation protocols (§2.2). We show that context can increase agreement between evaluators and substantially modify model win rates based on pairwise preference judgements (§4). Further, we show that context can enable us to gather more insights about model behavior, such as identifying contexts that align more closely with default model responses (§5) and assessing model sensitivity to different user contexts (§6). Our work provides a plug-and-play recipe for incorporating context into future language model evaluations.
2 Underspecification and Contexts
2.1 How Prevalent is Underspecification?
@@
2.2 How To Represent Context?
@@
In this work, we represent context as follow-up question-answer (QA) pairs, simulating an interactive scenario where an agent can query the user for additional details. For example, for a user issuing the query “Is coffee good for you?”, we can represent relevant medical context as “Q: Do you have hypertension? A: Yes”. In practical scenarios, this context could be collected through user interaction or be stored by the agent as memory from past conversations.
Desiderata for Questions
Formally, we require that each follow-up question is:
@@
Context typically describes user attributes or the scope of the user’s query (see Figure 2). User attributes include the user’s expertise, age group, location and other characteristics that affect the usefulness of the response. On the other hand, query scope describes the user’s preferences for the response, such as the specific topic or aspect they want covered, the desired length or format, or the need for references.
Desiderata for Answers
@@ 2.3 Synthetically Generating Contexts
Scalably collecting many, diverse contexts for study can be difficult. In this work, we demonstrate and evaluate the use of language models in generating follow-up questions and their answer sets.
@@ Results
Figure 3: Human validation of generated follow-up QA pairs shows that most follow-up questions are important to respond to the query and most answer sets are realistic, complete and diverse.
@@
To ensure that our contexts meet the criteria outlined in §2.2, we ask 251 human annotators recruited from Prolific to validate contexts for a random sample of 154 queries, where each example is annotated by 3 annotators. Each annotator is shown a query and the corresponding follow-up QAs, and asked to provide binary labels for the importance of each query, and the realisticness, completeness and diversity of the answer sets. Further annotation details are provided in Appendix A. The majority label percentages based on this validation are shown in Figure 3. We find that most follow-up questions ($\sim$76%) are found to be important, while answer sets are mostly complete ($\sim$75%), realistic ($\sim$90%) and diverse ($\sim$80%). This gives us confidence to rely on these generated contexts for our studies.
@@ 3 Designing Contextualized Evaluations
Model Pair NoCtxGen-CtxEval CtxGen-CtxEval
1 2 |Avg Diff.| 1 2 |Avg Diff.|
GPT-4o vs Gemini-1.5-Flash 5.29 5.14 1.33 7.69 7.40 1.01
Claude-3.5-Sonnet vs Llama-3.1-405B 5.40 5.29 1.64 8.03 7.61 0.96
Gemma-2-27B vs Jamba-1.5-Large 4.55 4.96 1.91 7.27 7.46 1.06
Table 3: Average number of constraints satisfied by different model pairs in context-aware evaluation settings. Shown here are the total number of follow-up QAs which are satisfied by candidate 1 and candidate 2 (in the order listed in the first column), as well as their absolute average difference. We note that the ability of models to follow instructions in the context varies.
Model Pair NoCtxGen-NoCtxEval NoCtxGen-CtxEval CtxGen-CtxEval
Agreement % w/ ties (w/o ties) $\kappa_{\text{Fleiss}}$ Agreement % w/ ties (w/o ties) $\kappa_{\text{Fleiss}}$ Agreement % w/ ties (w/o ties) $\kappa_{\text{Fleiss}}$
GPT-4o vs Gemini-1.5-Flash 68.03 (68.73) 0.2482 78.34 (78.40) 0.4672 74.51 (74.56) 0.3799
Claude-3.5-Sonnet vs Llama-3.1-405B 64.00 (65.04) 0.1657 74.57 (74.89) 0.3800 74.93 (75.13) 0.3837
Gemma-2-27B vs Jamba-1.5-Large 73.22 (74.40) 0.3726 80.98 (81.01) 0.5238 75.94 (75.99) 0.4165
Table 4: Autorater agreement statistics for different model pairs in all three evaluation settings. Note that agreement increases substantially in context-aware settings.
Model Pair NoCtxGen-NoCtxEval NoCtxGen-CtxEval CtxGen-CtxEval
1 2 Tie 1 2 Tie 1 2 Tie
GPT-4o vs Gemini-1.5-Flash 39.07 53.00 7.92 53.69 46.04 0.27 68.05 31.79 0.16
Claude-3.5-Sonnet vs Llama-3.1-405B 47.36 47.07 5.57 44.86 54.57 0.57 75.29 24.49 0.23
Gemma-2-27B vs Jamba-1.5-Large 55.16 39.16 5.68 46.65 52.91 0.44 63.89 35.93 0.17
Table 5: Win rates based on the majority autorater vote, for all model pairs and evaluation settings. Shown here are the win rates of candidate 1, candidate 2 (in the order listed in the first column) and their tie rate. Note that the preferred model (with a higher win rate) can vary across context-agnostic and context-aware settings.
@@ 3.1 Evaluation Settings
Figure 4: Evaluation protocols considered in our work.
@@
    Implicit Context Discovery (NoCtxGen-CtxEval): In this setting, responses are generated without context but evaluators are exposed to plausible context surrounding the query. This context is fixed across both responses for pairwise evaluation. Since models are left to assume contexts based on initial queries, we can use this setting to discover implicit biases in default model responses by providing context to the evaluators. We showcase an example of such an analysis in Section 5.
@@ 3.2 Model Pairs to be Evaluated
  • GPT-4o (rank 3) versus Gemini-1.5-Flash (rank 4), as top ranking proprietary models.
  • Claude-3.5-Sonnet (rank 5) versus Llama-3.1-405B (rank 6), chosen as a top ranking proprietary and open-weights model.
  • Gemma-2-27B (rank 19) versus Jamba-1.5-Large (rank 20), as similarly ranked open-weights contenders.
@@ 3.3 Collecting Human Evaluator Judgments
We recruit our human evaluators through Prolific and ask them to first complete a training task where they evaluate responses to queries in both context-agnostic and context-aware settings. In the context-agnostic setting, annotators are asked to give an overall preference for one of the two responses or indicate a Tie, along with a free-text justification. While in the context-aware setting, annotators are first asked to evaluate whether the constraints in each one of the follow-up QAs are satisfied by each of the two responses. Once they have provided these ratings, they are required to similarly provide an overall preference along with a free-text justification. The order of the two responses is always randomized. Each annotator is shown a single query and pairs of responses for one of the three evaluation settings, which is randomly chosen for the annotator. Every example is annotated by 3 different annotators and we collect ratings for 100 examples for all 3 model pairs and 3 settings, for a total of 100 * 3 * 3 * 3 = 2700 ratings. Further annotation details and interface screenshots are in Appendix A.
@@ 3.4 Collecting Model Evaluator Judgments
The use of language model-based evaluators (also known as autoraters) is becoming increasingly prevalent. Hence, we also use them to judge responses in all three settings. We use 5 autoraters for each candidate model pair, which includes the 6 models listed in §3.2 and Qwen2-72B-Instruct Yang et al. (2024), but always excluding the two candidate models to avoid self-preference bias Panickssery et al. (2024). In all three evaluation settings, autoraters are instructed to give an overall preference for one of the two responses or indicate a Tie, and then provide a free-text justification.
Model Pair NoCtxGen-NoCtxEval NoCtxGen-CtxEval CtxGen-CtxEval
Agreement % w/ ties (w/o ties) Agreement % w/ ties (w/o ties) Agreement % w/ ties (w/o ties)
GPT-4o vs Gemini-1.5-Flash 70.67 (73.33) 75.00 (77.15) 76.00 (76.36)
Claude-3.5-Sonnet vs Llama-3.1-405B 65.99 (67.75) 71.33 (72.84) 74.67 (77.19)
Gemma-2-27B vs Jamba-1.5-Large 71.33 (73.95) 74.00 (75.48) 75.33 (75.79)
Table 6: Human agreement statistics for different model pairs in all three evaluation settings. Note that the agreement is slightly higher for a subset of model pairs.
Model Pair NoCtxGen-NoCtxEval NoCtxGen-CtxEval CtxGen-CtxEval
1 2 Tie 1 2 Tie 1 2 Tie
GPT-4o vs Gemini-1.5-Flash 40.24 48.78 10.98 40.86 50.54 8.60 52.75 32.97 14.29
Claude-3.5-Sonnet vs Llama-3.1-405B 58.75 37.50 3.75 45.88 36.47 17.65 55.55 24.44 20.00
Gemma-2-27B vs Jamba-1.5-Large 38.82 54.12 7.06 36.67 50.00 13.33 49.43 34.48 16.09
Table 7: Win rates based on the majority human vote, for all model pairs and evaluation settings. Shown here are the win rates of candidate 1, candidate 2 (in the order listed in the first column) and their tie rate. Note that the preferred model (with a higher win rate) can vary across context-agnostic and context-aware settings.
Model Pair NoCtxGen-CtxEval CtxGen-CtxEval
Agreement % Agreement %
GPT-4o vs Gemini-1.5-Flash 82.42 81.40
Claude-3.5-Sonnet vs Llama-3.1-405B 79.43 77.33
Gemma-2-27B vs Jamba-1.5-Large 82.05 82.54
Table 8: Human agreement statistics for different model pairs in context-aware evaluation settings, when there was a difference of at least 1 in total follow-up QAs satisfied by the two responses. Note that agreement is higher for this subset of examples.

### v2

# Judging Language Model Responses to Underspecified Queries

@@ Abstract
Language model users often issue queries that lack specification, where the context under which a query was issued—such as the user’s identity, the query’s intent, and the criteria for a response to be useful—is not explicit. For instance, a good response to a subjective query like “What book should I read next?” would depend on the user’s preferences, and a good response to an open-ended query like “How do antibiotics work against bacteria?” would depend on the user’s expertise. This makes evaluation of responses to such queries an ill-posed task, as evaluators may make arbitrary judgments about the response quality. To remedy this, we present contextualized evaluations, a protocol that synthetically constructs context surrounding an underspecified query and provides it during evaluation. We find that the presence of context can 1) alter conclusions drawn from evaluation, even flipping benchmark rankings between model pairs, 2) nudge evaluators to make fewer judgments based on surface-level criteria, like style, and 3) provide new insights about model behavior across diverse contexts. Specifically, our procedure suggests a potential bias towards WEIRD (Western, Educated, Industrialized, Rich and Democratic) contexts in models’ “default” responses and we find that models are not equally sensitive to following different contexts, even when they are provided in prompts.111 Our code and data are available at https://www.github.com/allenai/ContextEval.
@@ 1 Introduction
Users of language models often issue queries that are underspecified Spärck-Jones et al. (2007); Clarke et al. (2009); Ziegler et al. (2019); Keyvan and Huang (2022); Herlihy et al. (2024), but common evaluation practices for language models do not account for this. Consider an evaluator presented with a language model’s response to an underspecified query such as “Is coffee good for you?” (Figure 1). A language model might respond with an explanation about benefits like antioxidants and mental alertness, but this output would be unacceptable to users with certain health conditions. Can an evaluator make an informed judgment about language model response quality without clear guidance on factors that reveal the intended user’s preferences, background or other necessary criteria for the response to be useful?
@@
In this work, we consider the role of context in the evaluation of language model responses to underspecified queries. We propose contextualized evaluations—a protocol to synthetically generate and incorporate diverse contexts (represented as follow-up question-answer pairs) when evaluating model responses to underspecified queries. By applying this procedure to queries from widely-used language model benchmark datasets, we investigate three main research questions:
First, we investigate whether providing context to evaluators has a substantial effect on the conclusions drawn from evaluation. We sample responses from pairs of language models and collect pairwise preference judgments from both model-based as well as human evaluators, in context-agnostic (only the query and model outputs) and context-aware settings (additionally with follow-up questions and answers to clarify the query). Our experiments show not only that inclusion of context during evaluation can significantly improve agreement between evaluators (3-10% absolute), but also that context-aware evaluation can even flip benchmark rankings between model pairs due to drastic changes in win rates. This raises concern for the reliability of findings produced from context-agnostic evaluations using today’s language model benchmarks, which we find to be full of underspecified queries (Table 1).
Finally, we investigate whether context helps us learn more about the ability of models to adapt to different user contexts. In our work, we use contexts to study biases exhibited by “default” model responses, i.e., those generated without context (§6), to underspecified queries. We observe patterns suggesting a potential bias where default model responses are better aligned with WEIRD (Western, Educated, Industrialized, Rich and Democratic) contexts Henrich et al. (2010), with weak to moderate agreement between human and model evaluators. Context can also allow us to directly evaluate the instruction-following and personalization abilities of models. We show this by evaluating sensitivity of models to different contextual attributes, where we find considerable disparity in the ability of models to adapt to different contexts, even when they’re provided as instructions in prompts7).
In summary, our findings suggest that underspecification can have a significant impact on the conclusions and insights drawn from evaluation. To address this, we propose contextualized evaluations, a simple and broadly applicable solution that involves synthesizing relevant context and injecting it into existing evaluation protocols (§3). We show that context can increase agreement between evaluators and substantially modify model win rates based on pairwise preference judgements (§5). Further, we show that context can enable us to gather more insights about model behavior, such as identifying contexts that align more closely with default model responses (§6) and assessing model sensitivity to different user contexts (§7). Our work provides a plug-and-play recipe for incorporating context into future language model evaluations.
2 How Prevalent is Underspecification?
@@
Table 1: Queries from five LM benchmark datasets (Chatbot Arena, AlpacaEval, MTBench, ExpertQA and KIWI), categorized based on the amount of underspecification. Queries can present multiple types (e.g., “What is the best team in the league?” is both incomplete and subjective).
@@
3 Representing Context as Follow-Up QA
Query Scope User Attributes
Focus & Angle Would you also like information on the causes of fever? Expertise & Familiarity What is your level of expertise in biology?
Style & Tone Do you want this explanation to be technical or high-level? Geographical Location Are you looking for global trends or trends in a specific region?
Intent Are you learning Flask for work or as a hobby? Age Group What age group do you belong to?
Level of Detail How detailed do you want the response to be? Language Fluency Would you like the response in French?
Format Do you want a step-by-step overview or just a summary? Interests What sports are you most interested in hearing about?
Need for Resources Would you like me to provide scientific studies and other references? Cultural Background What cultural perspective should be considered in the response?
Clarification Do you mean apple the company or fruit? Profession Are you asking as a researcher or developer?
Intended Audience Are you preparing this for a general audience or experts? Political Views Are you interested in the environmental perspective or economic impact?
Length How long would you like the summary to be? Gender Identity Should I focus on health issues among people of a specific gender?
Table 2: Taxonomy over types of common contextual attributes which can be lacking in underspecified queries, along with examples of follow-up questions for each attribute.
@@ 3 Representing Context as Follow-Up QA
We represent context as follow-up question-answer (QA) pairs, simulating an interactive scenario where an agent can seek clarifying details from the user before responding. For example, for a user issuing an underspecified query about coffee, we can represent relevant medical context:
User: “Is coffee good for you?
Agent: “Do you have hypertension?
User: “Yes.
where the follow-up QA pair (marked with ) serves as additional context.
In practical scenarios, this context could be collected through user interaction or be stored by the agent as memory from past conversations.
3.1 Desiderata for Questions
We require that each follow-up question is:
@@ 3.1 Desiderata for Questions
Context typically describes user attributes or the scope of the user’s query (see Table 2). User attributes include the user’s expertise, age group, location and other characteristics that affect the usefulness of the response. On the other hand, query scope describes the user’s preferences for the response, such as the specific topic or aspect they want covered, the desired length or format, or the need for references.
3.2 Desiderata for Answers
@@ 3.3 Synthetically Generating Contexts
Scalably collecting many, diverse contexts for study can be difficult. We demonstrate and evaluate the use of language models in generating follow-up questions and their answer sets.
@@ Results
Figure 2: Human validation shows that most generated follow-up questions are important for clarifying an underspecified query and most generated answer sets are realistic, complete and diverse.
@@
To ensure that our contexts meet the criteria outlined in §3, we ask 251 human annotators recruited from Prolific to validate contexts for a random sample of 154 underspecified queries, where each example is annotated by 3 annotators. Each annotator is shown the underspecified query and the corresponding follow-up QAs, and asked to provide binary labels for the importance of the follow-up question, and the realisticness, completeness and diversity of the answer sets. Further annotation details are provided in Appendix A. The majority label percentages based on this validation are shown in Figure 2. We find that most follow-up questions ($\sim$76%) are found to be important, while answer sets are mostly complete ($\sim$75%), realistic ($\sim$90%) and diverse ($\sim$80%). This gives us confidence to rely on these generated contexts for our studies.
@@ 4.1 Evaluation Settings
Figure 3: Our work defines two new evaluation settings—adaptive evaluation and implicit context discovery—distinctive from the standard evaluation paradigm which is context-agnostic.
@@
    Implicit Context Discovery (NoCtxGen-CtxEval): In this setting, responses are generated without context but evaluators are exposed to plausible context surrounding the query. This context is fixed across both responses for pairwise evaluation. Since models are left to assume contexts based on initial queries, we can use this setting to discover implicit biases in default model responses by providing context to the evaluators. We showcase an example of such an analysis in §6.
@@ 4.2 Model Pairs to be Evaluated
  • GPT-4o vs Gemini-1.5-Flash, which were Rank 3 vs Rank 4, selected as top-ranking proprietary models.
  • Claude-3.5-Sonnet vs Llama-3.1-405B, which were Rank 5 vs Rank 6, selected as similarly ranked proprietary and open-weights model.
  • Gemma-2-27B vs Jamba-1.5-Large, which were Rank 19 vs Rank 20, selected as similarly ranked open-weights contenders.
@@ 4.3 Collecting Human Evaluator Judgments
We recruit our human evaluators through Prolific and ask them to first complete a training task where they evaluate responses to queries in both context-agnostic and context-aware settings. In the context-agnostic setting, evaluators are asked to give an overall preference for one of the two responses or indicate a Tie, along with a free-text justification. While in the context-aware setting, evaluators are first asked to evaluate whether the constraints in each one of the follow-up QAs are satisfied by each of the two responses. Once they have provided these ratings, they are required to similarly provide an overall preference along with a free-text justification. The order of the two responses is always randomized. Each evaluator is shown a single query and pairs of responses for one of the three evaluation settings, which is randomly chosen for the evaluator. Every example is judged by 3 different evaluators and we collect ratings for 100 examples for all 3 model pairs and 3 settings, for a total of 100 * 3 * 3 * 3 = 2700 ratings. Further human evaluation details and interface screenshots are in Appendix A.
@@ 4.4 Collecting Model Evaluator Judgments
The use of language model-based evaluators (also known as autoraters or LM-as-judge) is becoming increasingly prevalent. Hence, we also use them to judge responses in all three settings. We use 5 autoraters for each candidate model pair, which includes the 6 models listed in §4.2 and Qwen2-72B-Instruct Yang et al. (2024), but always excluding the two candidate models to avoid self-preference bias Panickssery et al. (2024). In all three evaluation settings, autoraters are instructed to give an overall preference for one of the two responses or indicate a Tie, and then provide a free-text justification.
Evaluation Setting GPT-4o vs Gemini-1.5-Flash $\Delta$ Claude-3.5-Sonnet vs Llama-3.1-405B $\Delta$ Gemma-2-27B vs Jamba-1.5-Large $\Delta$
NoCtxGen-NoCtxEval 70.67 (73.33) - 65.99 (67.75) - 71.33 (73.95) -
NoCtxGen-CtxEval 75.00 (77.15) 4.33 (3.82) 71.33 (72.84) 5.34 (5.09) 74.00 (75.48) 2.67 (1.53)
CtxGen-CtxEval 76.00 (76.36) 5.33 (3.03) 74.67 (77.19) 8.68* (9.44)* 74.33 (75.79) 3.00 (1.84)
(a) Human Evaluator
Evaluation Setting GPT-4o vs Gemini-1.5-Flash $\Delta$ Claude-3.5-Sonnet vs Llama-3.1-405B $\Delta$ Gemma-2-27B vs Jamba-1.5-Large $\Delta$
NoCtxGen-NoCtxEval 68.03 (68.73) - 64.00 (65.04) - 73.22 (74.40) -
NoCtxGen-CtxEval 78.34 (78.40) 10.31* (9.67)* 74.57 (74.89) 10.57* (9.85)* 80.98 (81.01) 7.76* (6.61)*
CtxGen-CtxEval 74.51 (74.56) 6.48* (5.83)* 74.93 (75.13) 10.93* (10.09)* 75.94 (75.99) 2.72* (1.59)*
(b) Model Autorater
Table 4: Agreement rate (max 100%) across context-agnostic and context-aware evaluation settings. Pairs of model output judged for overall preference by three human evaluators (§4.3) and five autoraters (§4.4); agreement calculated both excluding ties and including ties (in parentheses). $\Delta$ shows absolute increase in agreement rate over baseline context-agnostic setting (NoCtxGen-NoCtxEval); significant differences (p $<$ 0.05) indicated with *.
Evaluation Setting GPT-4o vs Gemini-1.5-Flash $\Delta$ Claude-3.5-Sonnet vs Llama-3.1-405B $\Delta$ Gemma-2-27B vs Jamba-1.5-Large $\Delta$
NoCtxGen-NoCtxEval 68.03 - 64.00 - 73.22 -
NoCtxGen-CtxEval 82.42 14.39 79.43 15.43 82.05 8.83
CtxGen-CtxEval 81.40 13.37 77.33 13.33 82.54 9.32
Table 5: Agreement rates in human evaluator pairwise judgments (without ties) across context-aware evaluation settings, considering only those examples where the two responses differed by at least one in the number of follow-up QAs they satisfied. We find agreement is substantially higher for this subset of examples than the overall agreement rates reported in Table 4.
Evaluation Setting Human Evaluator Model Autorater
GPT-4o Gemini-1.5-Flash Tie GPT-4o Gemini-1.5-Flash Tie
NoCtxGen-NoCtxEval 40.24 48.78 10.98 39.07 53.00 7.92
NoCtxGen-CtxEval 40.86 50.54 8.60 53.69 46.04 0.27
CtxGen-CtxEval 52.75 32.97 14.29 68.05 31.79 0.16
Evaluation Setting Human Evaluator Model Autorater
Claude-3.5-Sonnet Llama-3.1-405B Tie Claude-3.5-Sonnet Llama-3.1-405B Tie
NoCtxGen-NoCtxEval 58.75 37.50 3.75 47.36 47.07 5.57
NoCtxGen-CtxEval 45.88 36.47 17.65 44.86 54.57 0.57
CtxGen-CtxEval 55.55 24.44 20.00 75.29 24.49 0.23
Evaluation Setting Human Evaluator Model Autorater
Gemma-2-27B Jamba-1.5-Large Tie Gemma-2-27B Jamba-1.5-Large Tie
NoCtxGen-NoCtxEval 38.82 54.12 7.06 55.16 39.16 5.68
NoCtxGen-CtxEval 36.67 50.00 13.33 46.65 52.91 0.44
CtxGen-CtxEval 49.43 34.48 16.09 63.89 35.93 0.17
Table 6: Model win rates across context-agnostic and context-aware evaluation settings. Pairs of model output judged for overall preference by three human evaluators (§4.3) and five autoraters (§4.4); wins are determined by majority vote. In each setting, the overall preferred model is indicated in bold.